# On the stability, complexity, and distribution of similarity classes of the longest edge bisection process for triangles

### v2

@@ 2.2. The geometric structure of the terminal quadruples
While running simulations, we have observed many remarkable phenomena regarding the geometric position of the points of $\mathrm{LEB}(z)$ in $D$. The following result, which describes the geometric distribution of the points of a terminal quadruple, is a forerunner of a complete geometric distribution analysis of $\mathrm{LEB}(z)$ (see §4.1).
@@
For every $z_{1}\in A$, the set $\{z_{1},z_{2},z_{3},z_{4}\}$ that is generated from $z_{1}$ lies on a hyperbolic circle centered at $\zeta=\frac{1}{2}+\frac{1}{2}i$. In particular, every terminal quadruple lies on an Euclidean circle centered at a point on the line $\mathrm{Re}(z)=\frac{1}{2}$.
@@
For $z\in\text{I}$, the point $L(z)$, is obtained by an inversion with respect to the geodesic $\left\{w\ \middle|\ |w|=\frac{\sqrt{2}}{2}\right\}$, that contains $\zeta=\frac{1}{2}+\frac{1}{2}i$. Since $\{z_{1},z_{2},z_{3},z_{4}\}\cap\text{I}\neq\emptyset$, we may assume that $z_{1}\in\text{I}$. The points $z_{1}$ and $L(z_{1})$ are both in region I and have the same hyperbolic distance from $\zeta$. Furthermore, for $z\in\text{I}$, the point $R(z)\in\text{IV}$ is obtained by a hyperbolic rotation of $\frac{\pi}{2}$ around $\zeta$. Thus $R(z_{1})$ and $RL(z_{1})$ are also at the same hyperbolic distance from $\zeta$ as $z_{1}$. ∎
@@ 2.3. Regions with $q(z)=1$
We begin with some observations regarding compositions of our basic functions $L$ and $R$ in certain regions.
@@
Recall that for $z\in\text{I}\cup\text{II}\cup\text{II}$ we have $R(z)=\frac{-1}{2z-2}$. One easily checks that $R$ maps the geodesics $\mathrm{Re}(z)=\frac{1}{2}$, $\left|{z-1}\right|=1$ and $\left|{z-\frac{1}{2}}\right|=\frac{1}{2}$ to the geodesics $\left|{z-\frac{1}{2}}\right|=\frac{1}{2}$, $\left|{z}\right|=\frac{1}{2}$ and $\mathrm{Re}(z)=\frac{1}{2}$, respectively, and hence maps the region $\text{I}\cup\text{II}\cup\text{III}$ to region IV. Since $L(z)=\frac{2z-1}{2z}$ and $R(z)=\frac{2\bar{z}-1}{2\bar{z}-2}$ for $z\in\text{IV}$, the assertion follows by a direct computation. ∎
@@
If $z\in\text{I}$
then $z$ belongs to the terminal quadruple $\mathrm{LEB}(z)=\{z,L(z),R(z),RR(z)\}$ and clearly $q(z)=1$.
If $z\in\text{II}$
@@
Since the triangles $z,R(z),RR(z)$ do not belong to a terminal quadruple and $L(z)\in\text{I}$, it follows that $q(z)=1$.
If $z\in\text{III}$
then using the same identities as in the previous case, (3) holds. As before, the triangles $z,R(z),RR(z)$ do not belong to a terminal quadruple. Thus, terminal quadruples can only arise from $\mathrm{LEB}(L(z))$. Notice that $L(z)\in\text{I}\cup\text{II}\cup\text{III}$. If $L(z)\in\text{I}\cup\text{II}$ then we are done, by the previous cases. Otherwise, $L(z)\in\text{III}$, and we may repeat the above argument. Namely, the singletons $\{L(z)\},\{RL(z)\}$ and $\{RRL(z)\}$, which do not belong to a terminal quadruple, are added to $\mathrm{LEB}(z)$, and terminal quadruples can only arise from $\mathrm{LEB}(LL(z))$. Since $\mathrm{LEB}(z)$ is finite, there exists some $k\in\mathbb{N}$ for which $L^{k}(z)\in\text{I}\cup\text{II}$, and we are done.
@@ 3. The bisection graph and its spectral properties
The following is a direct consequence of the Perron–Frobenius theorem (see [HJ13, Theorem 8.4.4]).
@@
Notice that with the above definition, even for non-generic points, a terminal quadruple corresponds to a directed $K_{2,2}$ subgraph of $\mathrm{G_{\mathrm{LEB}}}(z)$.
For a triangle $z$, we enumerate the similarity classes in $\mathrm{LEB}(z)$ by $[n]$, setting the similarity class of $z$ first, and consider the adjacency matrix $A\in M_{n}\left(\mathbb{R}\right)$ of $\mathrm{G_{\mathrm{LEB}}}(z)$ defined by (Note that our definition of the adjacency matrix is the transpose of the common definition.)
@@
    From the dynamics of the $\mathrm{LEB}$ process, we know that from every vertex there is a path to a vertex in a terminal quadruple. In particular, any strongly connected component $\alpha_{j}$ of $\mathrm{G_{\mathrm{LEB}}}(z)$ that does not correspond to a terminal quadruple has at least one outgoing edge. This implies that its corresponding adjacency matrix $B_{j}$ is irreducible (since $\alpha_{j}$ is strongly connected) and has at least one column sum that is strictly smaller than $2$. Applying Lemma 17, we deduce that any such component $\alpha_{j}$ has $\rho(\alpha_{j})<2$. Thus, the strongly connected components of $\mathrm{G_{\mathrm{LEB}}}(z)$ with spectral radius $2$ are precisely the terminal quadruples.
@@ 4.1. A remark on the geometric distribution of $\mathrm{LEB}(z)$
As we already mentioned in §2.2, we have observed several surprising properties of the geometric position of the points of $\mathrm{LEB}(z)$ in $D$, in particular, Proposition 8 characterized the geometric position of a terminal quadruple.
@@ 4.2. A remark on continuity and small perturbations
the $\varepsilon$-neighborhood of $F$. In [PP14, Lemma 3.2], the following non-increasing property of the metric $d$ was proven. For every $z,w\in D$ we have
@@
For every $z\in D$ we have the following: for every $\varepsilon>0$ and every $w\in D$ with $d(z,w)<\varepsilon$ we have $\mathrm{LEB}(w)\subseteq N_{\varepsilon}[\mathrm{LEB}(z)]$.

### v3

@@ 2.2. The geometric structure of the terminal quadruples
While running simulations, we have observed many remarkable phenomena regarding the geometric position of the points of $\mathrm{LEB}(z)$ in $\mathcal{D}$. The following result, which describes the geometric distribution of the points of a terminal quadruple, is a forerunner of a complete geometric distribution analysis of $\mathrm{LEB}(z)$ (see §4.1).
@@
For every $z_{1}\in\mathcal{A}$, the set $\{z_{1},z_{2},z_{3},z_{4}\}$ that is generated from $z_{1}$ lies on a hyperbolic circle centered at $\zeta=\frac{1}{2}+\frac{1}{2}i$. In particular, every terminal quadruple lies on an Euclidean circle centered at a point on the line $\mathrm{Re}(z)=\frac{1}{2}$.
@@
For $z\in\text{I}$, the point $L(z)$, is obtained by an inversion with respect to the geodesic $\left\{w\ \middle|\ |w|=\frac{\sqrt{2}}{2}\right\}$, that contains $\zeta=\frac{1}{2}+\frac{1}{2}i$. Since $\{z_{1},z_{2},z_{3},z_{4}\}\cap\text{I}\neq\emptyset$, we may assume that $z_{1}\in\text{I}$. The points $z_{1}$ and $L(z_{1})$ are both in region I and have the same hyperbolic distance from $\zeta$. Furthermore, for $z\in\text{I}$, the point $R(z)\in\text{IV}$ is obtained by a hyperbolic rotation of $\frac{\pi}{2}$ around $\zeta$. Thus $R(z_{1})$ and $RL(z_{1})$ are also at the same hyperbolic distance from $\zeta$ as $z_{1}$. The fact that a hyperbolic circle is a Euclidean circle is straightforward to verify using the two metrics. ∎
@@
2.3. The flow towards $\mathcal{A}$
In Proposition 11 below, we show that the orbit of every point eventually reaches $\mathcal{A}$. Since we did not find this fact explicitly in the literature, we provide a proof here. The proof relies on Lemma 10, showing that points get closer to $\mathcal{A}$ under repeated applications of $L$.
Lemma 10.
Let $d:\mathbb{H}\times\mathbb{H}\to[0,\infty)$ denote the hyperbolic metric. For every $z\in\mathcal{D}\setminus(\text{I}\cup\text{IV})$ we have $d(L(z),\zeta)<d(z,\zeta)$.
Proof.
Each piece $\tilde{L}$ of the piecewise function $L$ is an isometry of $\mathbb{H}$, hence for every $z\in\mathcal{D}$ we have $d(\tilde{L}(z),\zeta)=d(z,\tilde{L}^{-1}(\zeta))$. We proceed by case analysis for points $z$ in the different subregions of $\mathcal{D}$, where in each case we compute the perpendicular bisector of $\zeta$ and $\tilde{L}^{-1}(\zeta)$ and use it to deduce that $z$ is closer to $\tilde{L}^{-1}(\zeta)$.
  • If $z\in\text{II}\setminus\text{I}$, we have $\tilde{L}(z)=\frac{-1}{2z-1}$. Then $\tilde{L}^{-1}(\zeta)=\frac{\zeta-1}{2\zeta}=\frac{i}{2}$ and the perpendicular bisector of $\zeta$ and $\tilde{L}^{-1}(\zeta)$ is the geodesic $\mathrm{Re}(w)=\frac{1}{4}$.
  • If $z\in\text{III}\setminus\text{I}$, we have $\tilde{L}(z)=\frac{2\overline{z}}{2\overline{z}-1}$, and again $\tilde{L}^{-1}(\zeta)=\frac{\overline{\zeta}}{2\overline{\zeta}-2}=\frac{i}{2}$, and the perpendicular bisector is as in the previous case.
  • If $z\in\text{V}\setminus\text{I}$, we have $\tilde{L}(z)=2z$. Then $\tilde{L}^{-1}(\zeta)=\frac{\zeta}{2}$ and the perpendicular bisector is the geodesic $\left|{w}\right|=\frac{1}{2}$.
  • If $z\in\text{VI}\setminus\text{IV}$, we have $\tilde{L}(z)=1-2\overline{z}$. Then $\tilde{L}^{-1}(\zeta)=\frac{1-\overline{\zeta}}{2}=\frac{\zeta}{2}$, and we proceed as in the previous case.
In all of these cases, clearly $z$ is closer to $\tilde{L}^{-1}(\zeta)$ than to $z$, which completes the proof. ∎
Proposition 11.
For every $z\in\mathcal{D}$ there exists $n\in\mathbb{N}$ such that $L^{n}(z)\in\mathcal{A}$. In particular, $\mathrm{LEB}(z)\cap\mathcal{A}\neq\emptyset$.
Proof.
Let $z\in\mathcal{D}$ and assume for contradiction that $L^{n}(z)\notin\mathcal{A}$ for every $n\in\mathbb{N}$. Then by the FSC theorem, the set $\left\{L^{n}(z)\ \middle|\ n\in\mathbb{N}\right\}$ is a finite subset of $\mathcal{D}\setminus\mathcal{A}$. So there exists some $n_{0}\in\mathbb{N}$ for which $L^{n_{0}}(z)$ minimizes the distance from $\zeta$. But Lemma 10 asserts that $L^{n_{0}+1}(z)$ is closer. A contradiction. ∎
Another consequence is that $\mathcal{A}$ is exactly the locus of terminal quadruples.
Corollary 12.
For every $z\in\mathcal{D}$, if $\mathrm{LEB}(z)$ is a terminal quadruple then $z\in\mathcal{A}$.
Proof.
Suppose that $\mathrm{LEB}(z)$ is a terminal quadruple. By Proposition 9, the elements of $\mathrm{LEB}(z)$ are at the same distance from $\zeta$. For contradiction, assume that $z\in\mathcal{D}\setminus\mathcal{A}$, and consider two cases. If $z\notin\text{IV}$, by Lemma 10 we have $d(L(z),\zeta)<d(z,\zeta)$, a contradiction. If $z\in\text{IV}$, then $L(z)\in\text{II}$ and then $d(LL(z),\zeta)<d(z,\zeta)$, which is again a contradiction. ∎
@@ 2.4. Regions with $q(z)=1$
We begin with some observations regarding compositions of $L$ and $R$ in certain regions.
@@
Recall that for $z\in\text{I}\cup\text{II}\cup\text{III}$ we have $R(z)=\frac{-1}{2z-2}$. One easily checks that $R$ maps the geodesics $\mathrm{Re}(z)=\frac{1}{2}$, $\left|{z-1}\right|=1$ and $\left|{z-\frac{1}{2}}\right|=\frac{1}{2}$ to the geodesics $\left|{z-\frac{1}{2}}\right|=\frac{1}{2}$, $\left|{z}\right|=\frac{1}{2}$ and $\mathrm{Re}(z)=\frac{1}{2}$, respectively, and hence maps the region $\text{I}\cup\text{II}\cup\text{III}$ to region IV. Since $L(z)=\frac{2z-1}{2z}$ and $R(z)=\frac{2\bar{z}-1}{2\bar{z}-2}$ for $z\in\text{IV}$, the assertion follows by a direct computation. ∎
@@
Assume $z\in\text{I}$
Then $z$ belongs to the terminal quadruple
$\mathrm{LEB}(z)=\{z,L(z),R(z),RR(z)\}$
and clearly $q(z)=1$ (see e.g. [PP14, Lemma 4.4, Figure 8]).
Assume $z\in\text{II}$
@@
By Corollary 12, the triangles $z,R(z),RR(z)$ do not belong to a terminal quadruple and $L(z)\in\text{I}$, it follows that $q(z)=1$.
Assume $z\in\text{III}$
Then using the same identities as in the previous case, (3) holds. As before, the triangles $z,R(z),RR(z)$ do not belong to a terminal quadruple. Thus, terminal quadruples can only arise from $\mathrm{LEB}(L(z))$. Notice that $L(z)\in\text{I}\cup\text{II}\cup\text{III}$. If $L(z)\in\text{I}\cup\text{II}$ then we are done, by the previous cases. Otherwise, $L(z)\in\text{III}$, and we may repeat the above argument. Namely, the singletons $\{L(z)\},\{RL(z)\}$ and $\{RRL(z)\}$, which do not belong to a terminal quadruple, are added to $\mathrm{LEB}(z)$, and terminal quadruples can only arise from $\mathrm{LEB}(LL(z))$. By Proposition 11, there exists some $k\in\mathbb{N}$ for which $L^{k}(z)\in\mathcal{A}$, and we are done.
@@ 3. The bisection graph and its spectral properties
As a consequence of the Perron–Frobenius theorem (see [HJ13, Theorem 8.4.4]) we have:
@@
Notice that with the above definition, even for non-generic points, a terminal quadruple corresponds to a directed $K_{2,2}$ subgraph of $\mathrm{G_{\mathrm{LEB}}}(z)$ (this fact, in the generic case, already appeared in [RS75, Lemma 4]).
For a triangle $z$, we enumerate the similarity classes in $\mathrm{LEB}(z)$ by $[n]$, setting the similarity class of $z$ first, and consider the adjacency matrix $A\in M_{n}\left(\mathbb{R}\right)$ of $\mathrm{G_{\mathrm{LEB}}}(z)$ defined by
@@
Note that our definition of the adjacency matrix is the transpose of the common definition.
@@
    By Proposition 11, from every vertex there is a path to a vertex in a terminal quadruple. In particular, any strongly connected component $\alpha_{j}$ of $\mathrm{G_{\mathrm{LEB}}}(z)$ that does not correspond to a terminal quadruple has at least one outgoing edge. This implies that its corresponding adjacency matrix $B_{j}$ is irreducible (since $\alpha_{j}$ is strongly connected) and has at least one column sum that is strictly smaller than $2$. Applying Lemma 21, we deduce that any such component $\alpha_{j}$ has $\rho(\alpha_{j})<2$. Thus, the strongly connected components of $\mathrm{G_{\mathrm{LEB}}}(z)$ with spectral radius $2$ are precisely the terminal quadruples.
@@ 4.1. A remark on the geometric distribution of $\mathrm{LEB}(z)$
As we already mentioned in §2.2, we have observed several surprising properties of the geometric position of the points of $\mathrm{LEB}(z)$ in $\mathcal{D}$, in particular, Proposition 9 characterized the geometric position of a terminal quadruple.
@@ 4.2. A remark on continuity and small perturbations
the $\varepsilon$-neighborhood of $F$. In [PP14, Lemma 3.2], the following non-increasing property of the metric $d$ was proven. For every $z,w\in\mathcal{D}$ we have
@@
For every $z\in\mathcal{D}$ we have the following: for every $\varepsilon>0$ and every $w\in\mathcal{D}$ with $d(z,w)<\varepsilon$ we have $\mathrm{LEB}(w)\subseteq N_{\varepsilon}[\mathrm{LEB}(z)]$.